OpenStudy (anonymous):

Verify the identity: Use the fundamental , & sum-difference identities in the steps. ( sec y - tan y)² = 1-sin y/ 1+sin y

OpenStudy (anonymous):

\[(a-b)^2 = a^2 + b^2 - 2ab\] Use this and solve LHS..

OpenStudy (anonymous):

Solving or not??

OpenStudy (anonymous):

= sec² a + tan²a - 2 sec a tan a = 1+ tan²a + tan²a - 2 sec a tan a = 1+2tan²a + 2(1+tan a )(tan a) =1+ 4tan²a + 2 tan a .?

OpenStudy (anonymous):

You can do it by other method easily.. \[secy = \frac{1}{cosy}\] \[tany = \frac{siny}{cosy}\] So, \[(\frac{1}{cosy} - \frac{siny}{cosy})^2 = (\frac{1-siny}{cosy})^2 = \frac{(1-siny)^2}{\cos^2y}\] Open the parenthesis of numerator and then show me..

OpenStudy (anonymous):

See, what you have written on RHS I think there is some mistake.. Tell me clearly what we have to verify?

OpenStudy (anonymous):

verify that the 2 equations are equal/same

OpenStudy (anonymous):

NO..

OpenStudy (anonymous):

\[\left( \sec \alpha - \tan \alpha \right)^{2} = \left( 1-\sin \alpha / 1+\sin \alpha \right)\]

OpenStudy (anonymous):

Have you got it till here: \[\frac{(1-siny)^2}{\cos^2y}\]

OpenStudy (anonymous):

I don't understand how to get a minus sin y what I have is : = ( 1/ cos² y) + (sin² y/ cos ² y) - 2sec y tan y = ( 1+sin y/ cos y)² - 2 sec y tan y = (1+sin² y + 2sin y/ cos² y) - 2sec y tan y what happened to the ( 2sec y tan y ) in the one you did ?

OpenStudy (anonymous):

See tell me have you got till there??

OpenStudy (anonymous):

\[\cos^2x = 1 - \sin^2x = (1+sinx)(1-sinx)\] \[\frac{(1- siny)^2}{1 - \sin^2y} = \frac{ \cancel{1-siny} \times (1-siny)}{\cancel{(1-siny)} \times (1 + siny)} = \frac{1 - siny}{1 + siny}\]